OpenStudy (anonymous):

Whats the sum of { 5.5 + ( -5.5 + 2.3 ) without performing any additions? explain.

OpenStudy (aum):

Get rid of the inside parenthesis first: { 5.5 + ( -5.5 + 2.3 ) } = { 5.5 - 5.5 + 2.3 } = ?

OpenStudy (aum):

If there is a plus sign outside the parenthesis, then upon removing the parenthesis the signs of the inside quantities will remain the same as before. But if there is a minus sign outside the parenthesis, then upon removing the parenthesis, the signs of inside quantities will flip (that is + will change to - and - will change to +).

OpenStudy (aum):

Here there is a plus sign outside the inner parenthesis. So upon removing the parenthesis, the signs of the quantities inside the parenthesis will remain as before. Thus, { 5.5 + ( -5.5 + 2.3 ) } = { 5.5 - 5.5 + 2.3 } Now the +5.5 and the -5.5 will cancel each other out leaving you with just +2.3 or simply just 2.3